# Unified Source-Free Domain Adaptation

### v1

@@ 4.5 Source-Free Out-Of-Distribution Generalization
Data sets. This part evaluates LCFD in the SF-OODG setting based on four challenging ImageNet Variants dataset published in recent years. Specifically, IN-V2 (i.e., ImageNet-V2) [56] is an independent set consisting of natural images, collected from different sources, totally containing 10k images of 1000 ImageNet categories. IN-A (i.e., ImageNet-A) [57] is a challenging set consisting of these ”natural adversarial examples” (misclassified by a vanilla ResNet50 [58]), containing 7.5k images of 200 ImageNet categories. IN-R (i.e., ImageNet-R) [59] is an ImageNet sub-set with artistic renditions. It contains 30k images covering 200 ImageNet categories. IN-K (i.e., ImageNet-Sketch) [60] collects black and white sketches from the ImageNet validation set. This dataset contains 50k images from 1000 ImageNet categories.
Competitors. As shown in Tab. VII, besides these comparison methods adopted in the closed-set SFDA setting, we add four new comparisons, CoOP [44], CoCoOp [61], TPT [26] and ProGrad [62]. Among them, CoOP, CoCoOp, ProGrad perform prompt learning under the supervision of a few labels, whilst TPT proposes self-supervised prompt tuning via minimizing the multi-view entropy. Like the closed-set setting, the CLIP-based methods are marked with the backbone of image-encoder in CLIP.
Comparison results. Tab. VII reports the SF-OODG results on ImageNet Variations. In terms of average accuracy, our methods, LCFD-C-NR and LCFD-C-B32, are significantly ahead of the previous SFDA methods, improving by 10% at least. For CLIP and the CLIP-based methods, the advantage of LCFD is still noticeable. Taking results of the weak version as an example, while improving by 7.0% compared with CLIP-C-RN, LCFD-C-NR surpasses the second-best CLIP-based method TPT-RN by 3.8%.
TABLE VII: SF-OODG results (%) on ImageNet variants. L and M means label-required and multimodal, respectively.
@@
CLIP-RN ICML21 51.5 33.3 21.8 56.1 40.7
CLIP-B32 ICML21 54.8 40.8 29.5 66.2 47.8
CoOP-RN IJCV21 55.4 34.7 23.1 56.6 42.4
CoOP-B32 IJCV21 58.2 41.5 31.3 65.8 49.2
CoCoOp-RN CVPR22 55.7 34.5 23.3 57.7 42.8
CoCoOp-B32 CVPR22 56.6 40.7 30.3 64.1 47.9
TPT-RN NIPS22 54.7 35.1 26.7 59.1 43.9
ProGrad-RN ICCV23 54.7 34.4 23.1 56.8 42.2
ProGrad-B32 ICCV23 57.4 41.7 31.9 66.5 49.4
SHOT ICML20 62.5 38.4 2.1 42.7 36.4
GKD IROS21 62.3 38.2 2.1 42.2 36.2
NRC NIPS21 61.5 0.9 1.3 28.9 23.1
AdaCon CVPR22 50.9 18.5 2.0 38.6 27.5
CoWA ICML22 62.7 37.8 1.7 46.3 37.1
PLUE CVPR23 53.0 18.5 2.0 38.6 28.0
TPDS IJCV23 62.9 35.0 2.9 48.4 37.3
LCFD-C-RN 64.4 42.6 22.3 61.5 47.7
LCFD-C-B32 64.7 48.4 30.6 71.0 53.7
4.6 Further Discussion
During this entire discussion, all experiments are based on the strong version of LCFD, i.e., LCFD-C-B32. For brevity, we remove the suffix ’-C-B32’.
@@ 4.6.1 Analysis for Causal Factors Discovery
Figure 5: Accuracy changing of pseudo-label as training iteration going on.
Figure 5: Accuracy changing of pseudo-label as training iteration going on.
Figure 6: Stacked area chart of accuracy as Gaussian blur intensity enhancing.
Analysis of discovering process. In our LCFD framework, the latent external causal elements $S_{e}$ are captured and transformed into the observed pseudo-label. Thereby, the pseudo-label quality should be an indicator reflecting the discovery process. Here, we select the pseudo-label-based methods SHOT, COWA, GKD and CLIP’s zero shot (termed CLIP) as evaluation references. The comparison results are depicted in Fig. 6. Due to conducting $S_{e}$ discovery with the aid of CLIP, the pseudo-label accuracy of LCFD, i.e., LCFD-PL, is much better than previous methods as well as CLIP-PL. On the other hand, LCFD’s curve is on the top of LCFD-PL with a significant gap during the entire training process. This phenomenon shows that capturing internal causal elements $S_{i}$ is effective.
Causality evaluation of discovered causal factors. Ideally, when a model is designed to establish underlying causal relationships accurately, the model will exhibit immunity/ robustness to noise [27]. To evaluate the captured causal factors, we conduct a robustness analysis to noise. Specifically, we observe the accuracy change, whilst gradually increasing the kernel size of Gaussian blur (from 8 to 20) on the input target images. The comparison methods are CLIP, SHOT and LCFD-w-CLIP, which is a variation of LCFD where the pseudo-labels adopt the zero-shot results of CLIP directly.
As shown in Fig. 6, the performance of LCFD-w-CLIP, SHOT, CLIP deteriorates to varying degrees, as noise increases. In contrast, LCFD almost maintains the performance on all intensities. Specifically, compared with the accuracy at kernel size equals 8, LCFD-w-CLIP, SHOT, CLIP respectively decrease by 5.0%, 9.1% and 7.9% when kernel size equals 20, i.e., max noise intensity, whilst the decrease of LCFD is only 0.3%. The lower performance loss of LCFD implies the effectiveness of capturing causal factors by LCFD. In particular, together LCFD with CLIP, LCFD can work with impressive noise robustness as CLIP’s zero shot performance is sharply weakening. This result implies that LCFD’s performance is not due to CLIP’s excellent zero-shot capability, but deriving from the discovery of causal factors.
Also, comparing LCFD with LCFD-w-CLIP, the gap between them gradually broadens, implying capturing the external elements $S_{e}$ enables LCFD to gain immunity to noise. To gain a better understanding, we visualize the prediction distribution of LCFD-PL and CLIP at zero and max noise intensity, using the t-SNE tool. Fig. 7 displays the visualization results.
Figure 7: Prediction visualization of LCFD-PL, CLIP when Gaussian blur kernel sizes are 0 (N0) and 20 (N20), respectively. Their corresponding classification accuracy (%) is bracketed.
@@
When Gaussian blur kernel size is 0 (the first row), the clustering shape of LCFD-PL-N0 and CLIP-PL-N0 have a similar pattern: Samples with ambiguous semantics are mainly distributed in the middle area. As the kernel size reaches 20, CLIP-N20 continues the previous distribution pattern that ambiguous samples spread to a much wider central zone. Thereby, the noise drowns out the accurate category information provided by generated pseudo labels. As a result, the LCFD-w-CLIP’s performance deteriorates.
On the contrary, despite the accuracy being almost the same as CLIP-N20 (31%), LCFD-PL-N20 evolves a new pattern: While these ambiguous samples move to the periphery of the distribution, the remaining samples still cluster in the middle area with a shape similar to the LCFD-PL-N0 case. That is, as the noise enhancing, LCFD-PL-N20 maintains the distribution information. Correspondingly, with LCFD-PL-N20 guidance, LCFD achieves robustness to noise.
4.6.2 Model Analysis
Feature distribution visualization. Feature distribution is an intuitive evaluation means for classification. Here, we select five comparisons, including the source model (termed Source), CLIP, SHOT, TPDS and Oracle (trained on domain Cl with the real labels). Fig. 8 shows the logit visualization results in 3D Density chart. From Source to LCFD, the clustering effect gradually strengthens, and LCFD is the most similar to Oracle.
Ablation study. In the proposed framework, we discover the causal factors $S=\{S_{e},S_{i}\}$ in a separate way. We isolate the effect of this strategy by three variation LCFD methods. (i) LCFD-w-$S_{i}$: We only take $\cal{L}_{\rm{UN}}$ in Eq. (14) as the objective for model training, such that we test the effect of $S_{i}$ capturing. (ii) LCFD-w-$S_{e}$: We only perform the pseudo-label supervised learning, regulated by $\cal{L}_{\rm{PL}}$, where the pseudo-label is generated from $S_{e}$ discovery regulated by $\cal{L}_{\rm{EC}}$. Thus, the effect of $S_{e}$ capturing is tested. (iii) LCFD-w-CLIP: We directly adopt the zero-shot results of CLIP as the pseudo-labels regulating the $S_{i}$ discovery.
TABLE VIII: Ablation study results (%). CS, OS, PS and GS mean the closed-set, open-set, partial-set and Generalied SFDA settings, respectively.
@@ 4.6.2 Model Analysis
LCFD-w-$S_{i}$ 70.3 71.5 70.1 75.9 72.6 36.9
LCFD-w-$S_{e}$ 81.6 80.7 84.1 84.3 82.2 44.2
LCFD-w-CLIP 80.9 78.3 83.9 82.5 80.6 44.2
LCFD 83.3 83.4 85.8 84.5 83.7 48.4
TABLE IX: Scalability study (%) of IC instantiation.
@@
$\cal{L}_{\rm{UN}}$ LCFD+ENT 82.5 89.4 46.8
LCFD+NRC 83.0 87.9 47.5
LCFD+TPDS 83.5 90.2 45.0
@@
LCFD+${\rm{KL}}$ 83.5 89.1 49.1
LCFD 83.3 89.2 48.4
Tab. VIII reports the accuracy of these methods above in the four SFDA settings and SF-OODG. Specifically, when LCFD-w-$S_{i}$ or LCFD-w-$S_{e}$ works, the result outforms Source, but significantly lower than the full version LCFD. The results indicate that individual capturing of $S_{e}$ or $S_{i}$ is effective, and they jointly contribute to the final performance. Also, the comparison of LCFD-w-CLIP and LCFD confirm the effect of $\cal{L}_{\rm{EC}}$-based prompt learning for $S_{e}$ discovery.
Scalability study of IC instantiation. The objective $\cal{L}_{\rm{IC}}$ in Eq. (14) regulating $S_{i}$ discovery is a salable component with multiple implementations. To verify its scalability, this part respectively gives three other instantiations for the two regulators in $\cal{L}_{\rm{IC}}$. For $\cal{L}_{\rm{UN}}$, (i) the alone entropy regulating without category balance constraint (LCFD+ENT), (ii) data local structure-based regulating (LCFD+NRC) and (iii) domain shift control-based regulating (LCFD+TPDS) are selected as alternatives. For $\cal{L}_{\rm{PL}}$, the popular 1-Norm ($||\cdot||_{1}$), 2-Norm ($||\cdot||_{2}$) and KL divergence are employed.
As the results listed in Tab. IX, in the closed-set setting, LCFD+TPDS obtain the best results, and other methods have a gap of up to 2%. This indicates that these typical regulators do not substantially impact model performance in the closed-set setting. In the SF-OODG setting, besides LCFD+${\rm{KL}}$, all methods have a certain accuracy decrease compared with LCFD. For LCFD+ENT, removing category balance regulating leads to a decrease. The decrease of LCFD+NRC and LCFD+TPDS is due to the method not specifically considering the SF-OODG setting. Regarding the $\cal{L}_{\rm{PL}}$ group, the results indicate that probability distribution approximation (${\rm{KL}}$) outshines the features closing ($||\cdot||_{1}$, $||\cdot||_{2}$) in terms of maintaining the model performance caused by the external causal elements.
Effect of prompt initialization. In the LCFD method, we adopt prompt learning to encode the external causal factors $S_{e}$. In this part, we give four kinds of initialization strategies to figure out the effect of prompt initiation. As listed in Tab. X, except for conventional (i) constant strategy (row 1$\sim$2), (ii) sentence strategy (row 3$\sim$4) and (iii) phrase strategy (row 5$\sim$7), we introduce an interesting strategy of sentence initialization with uncertainty (row 8$\sim$9), considering the lacking of accurate supervision in SFDA.
Tab. X gives the comparison results of nine templates in the closed-set setting and SF-OODG. It can be observed that these templates do not lead to significant performance variance. These results indicate that the LCFD model is insensitive to prompt selection. Meanwhile, comparing the constant strategy and other strategies, the initialization with semantics is a better option, which also meets our expectations.
TABLE X: Prompt initialization study (%).
@@
1 ’X [CLASS].’ (#X=4) 82.6 89.1 48.4
2 ’X [CLASS].’ (#X=16) 81.5 88.5 46.1
3 ’There is a [CLASS].’ 83.6 89.2 48.6
4 ’This is a photo of a [CLS].’ 83.7 89.0 48.8
5 ’This is maybe a photo of a [CLS].’ 83.0 89.1 48.8
6 ’This is almost a photo of a [CLS].’ 83.7 89.2 48.4
7 ’This is definitely a photo of a [CLS].’ 83.5 88.9 48.1
8 ’a picture of a [CLS].’ 83.5 88.8 48.6
9 ’a photo of a [CLS].’ 83.3 89.2 48.4

### v4

@@ 4.6 Source-free out-of-distribution generalization
Datasets. We evaluate CausalDA in the SF-OODG setting on four challenging ImageNet Variants. Specifically, IN-V2 (i.e., ImageNet-V2) [43] is an independent set consisting of natural images, collected from different sources, a total of 10k images of 1000 ImageNet categories. IN-A (i.e., ImageNet-A) [18] is a challenging set consisting of these ”natural adversarial examples” (misclassified by a vanilla ResNet50 [16]), containing 7.5k images of 200 ImageNet categories. IN-R (i.e., ImageNet-R) [17] is an ImageNet sub-set with artistic renditions. It contains 30k images covering 200 ImageNet categories. IN-K (i.e., ImageNet-Sketch) [59] collects black and white sketches from the ImageNet validation set. This dataset contains 50k images from 1000 ImageNet categories.
Competitors. In addition to the methods compared in the Closed-set SFDA setting, we include four more: CoOP [70], CoCoOp [69], TPT [46], and ProGrad [71]. Among them, CoOP, CoCoOp, and ProGrad employ prompt learning with supervision from a few labeled samples, while TPT introduces self-supervised prompt tuning by minimizing multi-view entropy.
Comparison results. Tab. VIII reports the SF-OODG results. In average accuracy, our methods, CausalDA-C-NR and CausalDA-C-B32, outperform previous SFDA methods by a significant margin, achieving at least a 10% improvement. Even compared to CLIP and other CLIP-based methods, CausalDA maintains a clear advantage. For example, CausalDA-C-NR improves upon CLIP-C-RN by 7.0% and surpasses the second-best CLIP-based method, TPT-RN, by 3.8%. Among the “-B32” models, CausalDA-C-B32 excels with a margin of at least 10.5% over previous methods.
4.7 Further analysis
All experiments use the strong variant CausalDA-C-B32; the suffix “-C-B32” is omitted hereafter.
4.7.1 Causal vs. statistical SFDA
In this strategic evaluation, we contrast CausalDA with two state-of-the-art statistical SFDA methods, DIFO [53] and ProDe [52], both of which are equipped with the same ViL model. Our analysis focuses on two critical modeling aspects, examined across various SFDA settings: Domain fitness: Evaluated under Closed-set conditions; Domain generalization: Assessed across Generalized, Partial-set, Open-set, and SF-OODG settings.
Based on this evaluation, we draw the following key observations: (1) Domain fitness (Closed-set SFDA): As presented in Tab. IX, CausalDA is indeed outperformed by both DIFO and ProDe in the Closed-set setting. This is attributed to the inherent statistical learning nature of these alternatives, which tends to overfit their models towards the specific target domain. (2) Under the Generalized SFDA setting, CausalDA demonstrates superior performance in H-accuracy, also exhibiting the least drop in source domain accuracy (see S-accuracy). This is because the captured causality mitigates the risk of forgetting the source domain. (3) Except for the Closed-set scenario, CausalDA consistently stands out with strong performance on the other four settings. This highlights the superior domain generalization capabilities achieved through our method’s causal knowledge discovery. Importantly, our approach provides the best overall solution across all these diverse settings. Collectively, these results validate that our approach strikes the optimal balance between domain fitness and generalization, thereby laying a solid foundation to serve as a unified SFDA solution.
TABLE VIII: SF-OODG results (%) on ImageNet variants. L, ViL means label-required, ViL model-required, respectively.
@@ 4.7 Further analysis
CLIP-RN [42] ICML21 51.5 33.3 21.8 56.1 40.7
CLIP-B32 [42] ICML21 54.8 40.8 29.5 66.2 47.8
CoOP-RN [70] IJCV21 55.4 34.7 23.1 56.6 42.4
CoOP-B32 [70] IJCV21 58.2 41.5 31.3 65.8 49.2
CoCoOp-RN [69] CVPR22 55.7 34.5 23.3 57.7 42.8
CoCoOp-B32 [69] CVPR22 56.6 40.7 30.3 64.1 47.9
TPT-RN [46] NeurIPS22 54.7 35.1 26.7 59.1 43.9
ProGrad-RN [71] ICCV24 54.7 34.4 23.1 56.8 42.2
ProGrad-B32 [71] ICCV24 57.4 41.7 31.9 66.5 49.4
SHOT [31] ICML20 62.5 38.4 2.1 42.7 36.4
GKD [48] IROS21 62.3 38.2 2.1 42.2 36.2
NRC [64] NeurIPS21 61.5 0.9 1.3 28.9 23.1
AdaCon [5] CVPR22 50.9 18.5 2.0 38.6 27.5
CoWA [26] ICML22 62.7 37.8 1.7 46.3 37.1
PLUE [32] CVPR23 53.0 18.5 2.0 38.6 28.0
TPDS [49] IJCV24 62.9 35.0 2.9 48.4 37.3
SHOT-B32 [31] ICML20 69.4 43.6 1.6 58.2 43.2
GKD-B32 [48] IROS21 53.5 45.2 1.0 58.3 39.5
NRC-B32 [64] NeurIPS21 66.2 18.6 15.4 38.0 34.6
TPDS-B32 [49] IJCV24 69.0 38.9 0.6 58.7 41.8
CausalDA-C-RN 64.4 42.6 22.3 61.5 47.7
CausalDA-C-B32 64.7 48.4 30.6 71.0 53.7
TABLE IX: Causal vs. statistical SFDA. OH/VDA/DN126: Office-Home/VisDA/DomainNet-126. S/T: Source/Target domain; $H$: Harmonic mean accuracy. ${H}_{g}$: Average over all settings except closed-set.
Method Venue Closed-set Generalized Open-set Partial-set SF-OODG
OH VDA DN126 Avg. S T $H$ T T IN-V2 IN-K IN-A IN-R Avg. ${H}_{g}$}
Source 59.2 49.2 54.7 54.4 98.1 59.2 73.8 46.6 62.8 62.7 22.2 0.7 35.1 30.2 53.4
DIFO [53] CVPR24 83.1 90.3 80.0 84.5 78.0 83.1 80.5 75.9 85.6 59.6 37.7 25.2 66.4 47.2 72.3
ProDe [51] ICLR25 84.5 91.0 85.0 86.8 85.1 84.5 84.8 82.6 84.2 62.1 45.8 24.7 69.7 50.6 75.5
CausalDA 84.3 90.3 82.2 85.6 86.1 84.3 85.2 84.0 87.1 64.7 48.4 30.6 71.0 53.7 77.5
TABLE X: Unification comparison results (%). $H_{all}$: Average over all settings; $H_{wrg}$: The largest percentage shortfall of a method from the per-setting best across all settings; $H_{loso}$: Omit one setting at a time and average the scores over the rest.
Method $H_{wrg}$($\downarrow$) $H_{loso}$($\uparrow$) $H_{all}$($\uparrow$)
w/o Closed-set w/o Generalized setting w/o Open-set w/o Partial-set w/o SF-OODG
DIFO [53] 12.1 72.3 73.3 74.4 72.0 81.6 74.7
ProDe [51] 5.77 75.5 76.0 76.6 76.2 84.6 77.7
CausalDA 1.38 77.5 77.6 77.9 77.1 85.5 79.1
Unification comparison. To underscore CausalDA’s advantage over DIFO and ProDe, we perform a unification evaluation using the metrics in Eq. (18). The comparison in Tab. X shows that CausalDA attains the highest score in all scenarios. This indicates that CausalDA delivers robust, stable performance across diverse settings. Moreover, it exhibits no apparent weakness in any single scenario (cf. $H_{\mathrm{wrg}}$), and its all-round superiority is not driven by any individual setting (cf. $H_{\mathrm{loso}}$). Together, this evidence supports our claim of being towards a universally adaptable framework.
@@ 4.7.2 Analysis for causal factors discovery
We conduct causality analysis in two aspects: (1) complementarity of internal and external causality, and (2) analysis for the discovery process based on pseudo-label.
Complementarity of internal and external causality. As aforementioned, causal factors originate from both the domain data itself (encoded internal causality) and external knowledge sources (encoded external causality). To illustrate this complementarity, we conduct an intuitive empirical experiment on the target domain. Specifically, we fuse the zero-shot predictions of the source model (Source) and a ViL model like CLIP using a logical OR operation. That is, if either CLIP or Source correctly predicts the true category of an input instance $\boldsymbol{x}$, we consider $\boldsymbol{x}$ correctly classified. As shown in Fig. 6, Fusion-OR significantly outperforms both Source and CLIP individually, reinforcing the plausibility of our internal-external causality partition.
Pseudo-label-based analysis for discovery process. In CausalDA, the external causal elements $S_{e}$ are identified and converted into pseudo-labels. Thus, the quality of these pseudo-labels serves as an indicator of the discovery process. To evaluate this, we compare pseudo-label-based methods SHOT, COWA, and GKD, along with CLIP’s zero-shot predictions (denoted as CLIP). The results, shown in Fig. 7, demonstrate that CausalDA-PL achieves significantly higher pseudo-label accuracy than prior methods, benefiting from CLIP-assisted $S_{e}$ discovery. Furthermore, the performance curve of CausalDA remains consistently above that of CausalDA-PL, with a significant gap throughout training. This highlights the effectiveness of capturing internal causal elements $S_{i}$, further enhancing model performance.
Figure 6: Zero-shot fusion results (Fusion-OR) of CLIP and the source model (Source) under the Closed-set SFDA setting on the Office-Home, VisDA, and DomainNet-126 datasets. This fusion method follows the logical OR principle: An input instance is considered correctly classified if either CLIP or the source model predicts its ground truth category.
4.7.3 Model analysis
Feature distribution visualization. Feature distribution provides an intuitive way to evaluate classification performance. We compare with Source (source model), CLIP, SHOT, TPDS, and Oracle (trained on domain Cl with real labels). Fig. 8 presents the logit visualization results using a 3D density chart. As we move from Source to CausalDA, the clustering effect progressively improves, with CausalDA exhibiting the most structured distribution, closely resembling Oracle.
@@ 4.7.3 Model analysis
Ablation study. In CausalDA, we separately discover the causal factors $S\!=\!\{S_{e},S_{i}\}$. To isolate the effect of this strategy, we evaluate four variations: (i) CausalDA-w-$S_{i}$: We use only $\mathcal{L_{\rm{UN}}}$ from Eq. (17) as the objective for model training, allowing us to assess the impact of capturing $S_{i}$. (ii) CausalDA-w-$S_{e}$: We apply pseudo-label supervised learning, regulated by $\mathcal{L}_{\rm{SCE}}$, where the pseudo-labels are generated from $S_{e}$ discovery, regulated by $L_{\rm{EC}}$. This isolates the effect of capturing $S_{e}$. (iii) CausalDA-w-CLIP: We directly use CLIP’s zero-shot results as pseudo-labels to regulate the discovery of $S_{i}$. (iv) CausalDA-w-P1: We retain only the phase-1 objective, removing the loss $\mathcal{L_{\rm{IC}}}$ from Eq. (17).
Tab. XI presents the results of these methods across four SFDA settings and SF-OODG. Specifically, both CausalDA-w-$S_{i}$ and CausalDA-w-$S_{e}$ outperform the Source model, but fall significantly behind the full CausalDA version. This suggests that capturing either $S_{e}$ or $S_{i}$ independently is effective, but both are necessary for optimal performance. The comparison between CausalDA-w-CLIP and CausalDA highlights the importance of $L_{\rm{EC}}$-based prompt learning for $S_{e}$ discovery. Finally, CausalDA-w-P1 performs considerably worse than CausalDA, emphasizing the value of our two-stage design in the causal factor capture pipeline.
@@
Scalability study of IC instantiation. The objective $\mathcal{L}_{\rm{IC}}$ in Eq. (17), which regulates $S_{i}$ discovery, is a component with multiple implementation options. To evaluate its scalability, we present three alternative instantiations for the two regulators in $\mathcal{L}_{\rm{IC}}$. For $\mathcal{L}_{\rm{UN}}$, we consider the following alternatives: (i) CausalDA+ENT: Entropy regularization without category balance constraints, (ii) CausalDA+NRC: Data local structure-based regularization, (iii) CausalDA+TPDS: Domain shift control-based regularization. For $\mathcal{L}_{\rm{SCE}}$, we test three popular methods: 1-Norm ($||\cdot||_{1}$), 2-Norm ($||\cdot||_{2}$), and KL Divergence.
As shown in Tab. XII, in the Closed-set setting, CausalDA+TPDS achieves the best results, with other methods lagging by up to 1.9% (on VisDA). This indicates that these typical regulators do not significantly affect model performance in this setting. In the SF-OODG setting, all methods (except CausalDA+KL) show a slight accuracy drop compared to CausalDA. For CausalDA+ENT, removing the category balance regularization leads to a performance decrease. The performance loss for CausalDA+NRC and CausalDA+TPDS is attributed to the fact that these methods do not specifically address the SF-OODG setting.
For $\mathcal{L}_{\rm{SCE}}$, the results show that probability distribution approximation (KL Divergence) outperforms feature-based methods ($||\cdot||_{1}$ and $||\cdot||_{2}$) in preserving model performance in the presence of external causal elements.
TABLE XI: Ablation study results (%). CS, OS, PS, and GS mean the Closed-set, Open-set, Partial-set, and Generalized SFDA settings, respectively.
@@
CausalDA-w-$S_{i}$ 71.3 72.1 71.6 76.6 72.9 36.9
CausalDA-w-$S_{e}$ 82.5 81.4 85.5 84.9 83.6 44.2
CausalDA-w-CLIP 82.0 78.8 85.1 83.3 82.3 44.2
CausalDA-w-P1 80.3 38.1 80.7 78.1 69.3 40.8
CausalDA 84.3 84.0 87.1 85.2 85.1 48.4
TABLE XII: Scalability study (%) of IC instantiation.
@@
$\cal{L}_{\rm{UN}}$ CausalDA+ENT 83.5 90.4 46.8
CausalDA+NRC 83.9 89.0 47.5
CausalDA+TPDS 84.6 90.9 45.0
@@
[…]
CausalDA 84.3 90.3 48.4
TABLE XIII: Prompt initialization study (%).
@@
1 ’X [CLASS].’ (#X=4) 83.6 90.1 48.4
2 ’X [CLASS].’ (#X=16) 82.5 89.6 46.1
3 ’There is a [CLASS].’ 84.6 90.2 48.6
4 ’This is a photo of a [CLASS].’ 84.7 89.9 48.8
5 ’This is maybe a photo of a [CLASS].’ 83.9 90.2 48.8
6 ’This is almost a photo of a [CLASS].’ 84.8 90.3 48.4
7 ’This is definitely a photo of a [CLASS].’ 84.5 90.0 48.1
8 ’a picture of a [CLASS].’ 84.6 89.8 48.6
9 ’a photo of a [CLASS].’ 84.3 90.3 48.4
@@
Effect of prompt initialization. In CausalDA, we employ prompt learning to encode the external causal factors $S_{e}$. Here, we examine four different prompt initialization strategies (Tab. XIII): (i) the conventional constant strategy (rows 1–2), (ii) the sentence strategy (rows 3–4), (iii) the uncertainty strategy (rows 5–7), and (iv) an innovative approach of sentence initialization with different phrase (rows 8–9), which considers the lack of precise supervision in SFDA.
Tab. XIII shows the comparison results for these nine templates across both the Closed-set setting and SF-OODG. The results indicate that the different templates do not cause significant performance differences, suggesting that CausalDA is relatively insensitive to prompt selection. However, when comparing the constant strategy to the others, it becomes clear that initializing with semantic information is the more effective option, which aligns with our expectations.